an object is rotating on a circular path at 4 revolutions per minute. the linear velocity of the object is 400 feet per minute. what is the radius of the circle and what is the angular velocity of the rotating platform?
1
Expert's answer
2011-11-24T09:30:30-0500
T=60/4=15 -- period
n=1/T=0.0666 -- frequency
L=400/4=100 --length of range
L=2*pi*R
R=L/(2*pi)=100/6.28=15.92 -- radius
w=2*pi*0.0666=6.28*0.066=4.18248 -- angular velocity